### v1

# White-box Multimodal Jailbreaks Against Large Vision-Language Models

Abstract
Recent advancements in Large Vision-Language Models (VLMs) have underscored their superiority in various multimodal tasks. However, the adversarial robustness of VLMs has not been fully explored. Existing methods mainly assess robustness through unimodal adversarial attacks that perturb images, while assuming inherent resilience against text-based attacks. Different from existing attacks, in this work we propose a more comprehensive strategy that jointly attacks both text and image modalities to exploit a broader spectrum of vulnerability within VLMs. Specifically, we propose a dual optimization objective aimed at guiding the model to generate affirmative responses with high toxicity. Our attack method begins by optimizing an adversarial image prefix from random noise to generate diverse harmful responses in the absence of text input, thus imbuing the image with toxic semantics. Subsequently, an adversarial text suffix is integrated and co-optimized with the adversarial image prefix to maximize the probability of eliciting affirmative responses to various harmful instructions. The discovered adversarial image prefix and text suffix are collectively denoted as a Universal Master Key (UMK). When integrated into various malicious queries, UMK can circumvent the alignment defenses of VLMs and lead to the generation of objectionable content, known as jailbreaks. The experimental results demonstrate that our universal attack strategy can effectively jailbreak MiniGPT-4 with a 96% success rate, highlighting the vulnerability of VLMs and the urgent need for new alignment strategies.
Disclaimer: This paper may contain content that is potentially disturbing and offensive.
1 Introduction
The recent advancements in Large Vision-Language Models (VLMs) such as OpenAI’s GPT-4 (Achiam et al., 2023) and Google’s Flamingo (Alayrac et al., 2022) have received broad attention from academics, industry, and the general public. These developments have also garnered widespread attention from AI safety researchers, who are increasingly focused on assessing the adversarial robustness of multimodal models. While VLMs are expected to exhibit the same alignment characteristics as Large Language Models (LLMs), providing helpful responses to user queries while rejecting requests that may cause harm, the integration of additional visual modality to base language models introduces novel vulnerabilities. Several methods (Carlini et al., 2024; Bagdasaryan et al., 2023; Bailey et al., 2023; Qi et al., 2023; Shayegani et al., 2023a) have successfully manipulated VLMs to produce harmful content by launching single-modality attacks on the continuous and high-dimensional visual modality. Conversely, research on purely text-based attacks is limited, attributed to a prevailing consensus regarding their ineffectiveness in breaching the defense mechanisms of well-aligned VLMs, due to the discrete and lower-dimensional nature of text (Carlini et al., 2024).
Figure 1: Example of Jailbreak Attack on MiniGPT-4 (Zhu et al., 2023). The proposed Universal Master key (UMK) helps arbitrary harmful queries bypass alignment constraints.
In contrast, this paper presents an initial effort in introducing a text-image multimodal attack strategy, aiming to uncover a wide range of intrinsic vulnerabilities within VLMs. Furthermore, we have observed that while the previous unimodal universal attack strategy (Qi et al., 2023) increases response toxicity by perturbing images in the absence of text input, it also undermines the model’s adherence to instructions. Therefore, we propose a novel dual optimization objective strategy to address this issue.
Specifically, we employ white-box attacks to identify a multimodal adversarial example capable of jailbreaking VLMs, which we refer to as the Universal Master Key (UMK). The proposed UMK comprises an adversarial image prefix and an adversarial text suffix. We expect that by attaching the UMK to arbitrary malicious user queries, it will circumvent the alignment defenses of VLMs and provoke the generation of objectionable content. To discover such a UMK, we start by initializing an adversarial image prefix from random noise and optimizing it without text input to maximize the model’s probability in generating harmful content, thereby imbuing the adversarial image prefix with toxic semantics. Subsequently, we introduce an adversarial text suffix for joint optimization with the image prefix to maximize the probability of affirmative model responses. This strategy is based on the intuition that encouraging the model to produce affirmative responses may increase its tendency to engage in undesired behaviors, rather than rejecting responses. Figure 1 exemplifies our proposed method’s efficacy in attacking MiniGPT-4 (Zhu et al., 2023).
The main contributions of our work can be summarized as:
  • To the best of our knowledge, we are the first to introduce text-image multimodal adversarial attack against VLMs, systematically exploiting the vulnerabilities inherent in these models.
  • We propose a multimodal attack strategy with dual optimization objectives. We first enhance image toxicity by optimizing the adversarial image prefix, and then jointly optimize both the adversarial image prefix and adversarial text suffix to maximize the probability of affirmative model responses.
  • Extensive experiments on benchmark datasets demonstrate that the proposed Universal Master Key (UMK) can universally jailbreak the VLMs with a remarkable success rate, surpassing existing unimodal methods.
Figure 2: Overview of our multimodal attack strategy: The Universal Master Key (UMK) comprises an adversarial image prefix $X^{p}_{adv}$ and an adversarial text suffix $X^{s}_{adv}$. We first optimize $X^{p}_{adv}$ to maximize the generation probability of harmful content without text input to infuse toxic semantics. Subsequently, we concatenate the malicious user query with $X^{s}_{adv}$, and jointly optimize $X^{p}_{adv}$ and $X^{s}_{adv}$ to maximize the generation probability of affirmative responses, e.g., ‘Sure, here’s instruction for doing ******** (a bad thing)’.
2 Related Work
2.1 Large Vision-Language Models
Large Vision-Language Models (VLMs) are vision-integrated Large Language Models (LLMs) that receive input in text and image formats and generate free-form text output for multimodal tasks. VLMs typically leverage pre-trained LLMs and image encoders, connected by a text-image feature alignment module, enabling the language model to comprehend image features and engage in deeper question-answering reasoning. Taking several open-source VLMs as examples, LLaVA (Liu et al., 2024) leverages the language-only GPT-4 (Achiam et al., 2023) to generate multimodal language-image instruction-following data, enhancing zero-shot capabilities on new tasks through instruction tuning. It connects the open-source visual encoder CLIP (Radford et al., 2021) with the language decoder LLaMA (Touvron et al., 2023), and performs end-to-end fine-tuning on the generated visual-language instruction data. MiniGPT-4 (Zhu et al., 2023) attributes the advanced multimodal generation capabilities of GPT-4 (Achiam et al., 2023) to its integration of more sophisticated LLMs. To achieve similar capabilities, it employs a single linear projection layer to align the pretrained ViT (Dosovitskiy et al., 2020) and Q-Former (Li et al., 2023) with a frozen Vicuna (Chiang et al., 2023). The model is first pre-trained on an extensive dataset of aligned image-text pairs to acquire foundational visual-language knowledge. Subsequently, it undergoes fine-tuning using a smaller, higher-quality dataset of image-text pairs. In addition, MiniGPT-4 carefully designs dialogue templates to enhance the model’s generative reliability and usability. InstructBLIP (Dai et al., 2023) undertakes a rigorous and extensive analysis of vision-language instruction tuning, leveraging pre-trained BLIP-2 (Li et al., 2023) models. The study compiles 26 publicly accessible datasets and reformats them for instruction tuning. Moreover, InstructBLIP introduces an innovative instruction-aware Query Transformer. This component is designed to extract informative features specifically aligned with the provided instructions, enhancing the model’s capability to interpret and respond to instruction-based queries. Despite the exciting potential demonstrated by VLMs, the incorporation of an additional modality inadvertently introduces more vulnerabilities, thereby creating previously non-existent attack surfaces (Shayegani et al., 2023b). In our work, we evaluate the robustness of VLMs against the proposed multimodal attack, revealing an alarming attack success rate of 96% on MiniGPT-4 (Zhu et al., 2023), emphasizing the urgent need for new alignment strategies to rectify this critical vulnerability.
2.2 Attacks Against Multimodal Models
To attack multimodal models, Greshake et al. (Greshake et al., 2023) explored the effectiveness of manually injecting deceptive text into input images. In contrast, Other studies have proposed more sophisticated image-domain adversarial attack methods. Carlini et al. (Carlini et al., 2024) fixed the beginning portion of toxic target output and optimized the input images to increase its likelihood. Bagdasaryan et al. (Bagdasaryan et al., 2023) and Bailey et al. (Bailey et al., 2023) employed a similar strategy by using teacher-forcing techniques to generate the attacker-chosen text that may not be directly related to toxic content. Another white-box attack proposed by Qi et al. (Qi et al., 2023) adopts principles similar to Bagdasaryan et al. (Bagdasaryan et al., 2023), aiming to find an universal adversarial visual input. Specifically, the attack no longer focuses on specific output sentences but tries to maximize the probability of generating derogatory output from a corpus containing 66 harmful sample sentences. This strategy is inspired by Wallace et al. (Wallace et al., 2019), who also utilized optimization algorithms based on discrete search (Ebrahimi et al., 2017) to find universal adversarial triggers in token space, increasing the probability of generating a small group of harmful sentences. Since Qi et al. (Qi et al., 2023) did not provide a name for their method, we designate it as the Visual Adversarial Jailbreak Method (VAJM) for easier reference in subsequent sections. Shayegani et al. (Shayegani et al., 2023a) proposed attacking publicly available visual encoders such as CLIP (Radford et al., 2021) used in the multimodal models, thereby eliminating the need for complete white-box access. While the aforementioned approaches have demonstrated impressive results, they mainly focus on exploring the adversarial robustness under the unimodal attacks, based on the consensus that attacking in the continuous image space is more effective than attacking in the discrete token space. However, this focus may lead to underutilization of the full attack surfaces available in multimodal models. In contrast, our proposed method explores a broader range of vulnerabilities inherent in VLMs through attacks on text-image multimodalities. Moreover, the universal adversarial attack method proposed by Qi et al. (Qi et al., 2023) has been found to compromise the model’s adherence to user instructions. In comparison, our method employs dual optimization objectives aimed at guiding the model to generate affirmative responses with high toxicity, effectively alleviating this issue.
3 Methodology
3.1 Threat Model
Attack Goals. We consider single-turn conversations between a malicious user and a VLM chatbot. The attacker attempts to trigger a number of harmful behaviors by circumventing VLM’s security mechanisms, e.g., generating unethical content or dangerous instructions restricted by system prompts or RHLF alignment technique. Adversary Capabilities. We assume that the attacker has white-box access to the target VLM, a reasonable assumption given the advanced capabilities and extensive pre-training knowledge inherent in state-of-the-art open-source VLMs. These models are sufficient to provide attackers with the insights necessary to generate malicious content, such as detailed instructions for making a bomb or materials that promote gender discrimination.
3.2 Proposed Attack
3.2.1 Formalization
For simplicity, we omit the implementation details of converting raw input to feature embeddings. Consider a VLM $f_{\theta}$ as a mapping function from an image input $X_{i}$ and a text input $X_{t}$ to a probability distribution over the text output $Y$:
$p(Y|X_{i},\,X_{t})=f_{\theta}([X_{i},\,X_{t}])$
The proposed Universal Master Key (UMK) comprises an adversarial image prefix $X^{p}_{adv}$ and an adversarial text suffix $X^{s}_{adv}$. To jailbreak VLM $f_{\theta}$ and elicit harmful behavior $Y_{harm}$ with malicious user query $X_{harm}$, the attack is constructed as:
$p(Y_{harm}|X^{p}_{adv},\,X_{harm}||X^{s}_{adv})=f_{\theta}([X^{p}_{adv},\,X_{harm}||X^{s}_{adv}])$
where the adversarial image prefix $X^{p}_{adv}$ serves as the image input, while the malicious user query $X_{harm}$ is concatenated with the adversarial text suffix $X^{s}_{adv}$ to form the text input. ‘||||| |’ denotes string concatenation.
3.2.2 Methodology Intuition
VAJM (Qi et al., 2023) generates universal adversarial images by guiding the model to produce harmful content in the absence of text input, but this approach undermines the model’s ability to faithfully follow user instructions. In the development of jailbreak attacks against Large Language Models (LLMs), it is common practice to induce the model to respond to user queries in an affirmative manner, prioritizing task completion over query rejection. In manual attacks, this often involves prompting the model to begin its response with ‘sure’. Similarly, in Greedy Coordinate Gradient (GCG) (Zou et al., 2023), the optimization objective is to induce the model to affirmatively repeat user queries. However, we also discover that affirmatively repeating user queries does not necessarily lead to the generation of harmful content. The models still face conflicting objectives between task completion and benign alignment. Inspired by VAJM (Qi et al., 2023) and GCG (Zou et al., 2023), we propose a dual optimization objective that combines the advantages of both methods to maximize the probability of the model generating affirmative responses with high toxicity. Specifically, we begin by compromising the model’s ethical alignment through the infusion of toxic semantics into the adversarial image prefix. Subsequently, we jointly optimize the adversarial image prefix and adversarial text suffix to encourage the model to generate affirmative responses, with the expectation that the model will accomplish tasks in a state conducive to producing harmful content.
3.2.3 Embedding Toxic Semantics into Adversarial Image Prefix
Inspired by (Qi et al., 2023), we establish a corpus containing several few-shot examples of harmful sentences $S:=\{s_{i}\}_{i=1}^{m}$. The methodology for embedding toxic semantics into the adversarial image prefix $X^{p}_{adv}$ is straightforward: we initialize $X^{p}_{adv}$ with random noise and optimize it to maximize the generation probability of this few-shot corpus in the absence of text input. Our optimization objective is formulated as follows:
$X^{p}_{adv}:=\underset{X^{p}_{adv}}{\operatorname{argmin}}\sum_{i=1}^{m}-\log\left(p(s_{i}\mid X^{p}_{adv},\,\emptyset)\right)$
where $\emptyset$ denotes an empty text input. This optimization problem can be efficiently solved using prevalent techniques in image adversarial attacks, such as Projected Gradient Descent (PGD) (Madry et al., 2017).
3.2.4 Text-Image Multimodal Optimization for Maximizing Affirmative Response Probability
To maximize the probability of generating affirmative responses, we further introduce an adversarial text suffix $X^{s}_{adv}$ in conjunction with the adversarial image prefix $X^{p}_{adv}$, which is imbued with toxic semantics. This multimodal attack strategy aims to thoroughly exploit the inherent vulnerabilities of VLMs. Specifically, our optimization objective is as follows:
$X^{p}_{adv},\,X^{s}_{adv}:=\underset{X^{p}_{adv},\,X^{s}_{adv}}{\operatorname{argmin}}\sum_{i=1}^{n}-\log\left(p(t_{i}\mid X^{p}_{adv},\,g_{i}||X^{s}_{adv})\right)$
A few-shot corpus composed of goal-target pairs $\{g_{i},t_{i}\}_{i=1}^{n}$ is employed. Here, the goals $\{g_{i}\}_{i=1}^{n}$ represent the malicious user queries, while the targets $\{t_{i}\}_{i=1}^{n}$ are the corresponding affirmative repetitions of these queries, each prefixed with ‘Sure, here is’.
The shared feature space between the image embeddings and the text token embeddings in VLMs facilitates concurrent updates of both $X^{p}_{adv}$ and $X^{s}_{adv}$ through a single backpropagation pass. This mechanism is analogous to the simultaneous update of distinct pixels in a single image in conventional image adversarial attack. We employ Projected Gradient Descent (PGD) (Madry et al., 2017) to update the adversarial image prefix $X^{p}_{adv}$, and Greedy Coordinate Gradient (GCG) (Zou et al., 2023) to update the adversarial text suffix $X^{s}_{adv}$. GCG is the current state-of-the-art text-based attack strategy against LLMs. It utilizes the gradients with respect to the one-hot token indicators to identify a set of promising replacement candidates at each token position, and then chooses the substitution that results in the maximum loss reduction via a forward pass.
The overview of our attack is illustrated in Figure 2, while the detailed strategy is summarized in Algorithm 1.
Algorithm 1 Multimodal Attack Strategy with Dual Optimization Objectives
0:  VLM model $f_{\theta}$, harmful sentences corpus $S:=\{s_{i}\}_{i=1}^{m}$, goal-target pairs corpus $D:=\{g_{i},t_{i}\}_{i=1}^{n}$, image-text attack ratio $r$, batch size $b$, step size $\alpha$, number of candidates $n$, iteration counts $N_{1},N_{2}$;
1:  Initialize $X^{p}_{adv}$ with random noise;
2:  for $k=1,...,N_{1}$ do
3:     Select $b$ samples from the corpus $S$, forming $S_{k}:=\{s^{\prime}_{i}\}_{i=1}^{b}$;
4:     $\mathcal{L}(X^{p}_{adv},S_{k})=\sum_{i=1}^{b}-\log\left(p(s^{\prime}_{i}\mid X^{p}_{adv},\,\emptyset)\right)$;
5:     $X^{p}_{adv}=\operatorname{clip}\left(X^{p}_{adv}+\alpha\operatorname{sign}\left(\nabla_{x^{p}}\mathcal{L}(X^{p}_{adv},S_{k})\right)\right)$;
6:  end for
7:  Initialize $X^{s}_{adv}$ using a predefined strategy;
8:  for $j=1,...,N_{2}$ do
9:     Select $b$ samples from the corpus $D$, forming $D_{j}:=\{g^{\prime}_{i},t^{\prime}_{i}\}_{i=1}^{b}$;
10:     $\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})=\sum_{i=1}^{b}-\log\left(p(t^{\prime}_{i}\mid X^{p}_{adv},\,g^{\prime}_{i}||X^{s}_{adv})\right)$;
11:     $X^{p}_{adv}=\operatorname{clip}\left(X^{p}_{adv}+\alpha\operatorname{sign}\left(\nabla_{x^{p}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})\right)\right)$;
12:     if $j\%r==0$ then
13:        Compute coordinate gradient $\nabla_{x^{s}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})$;
14:        Obtain a set of candidate suffixes with single-token substitution $\{X^{s}_{adv}\}^{n}$;
15:        $X^{s}_{adv}=\underset{\{X^{s}_{adv}\}^{n}}{\operatorname{argmin}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})$;
16:     end if
17:  end for
18:  Obtain the Universal Master Key (UMK) composed of $X^{p}_{adv}$ and $X^{s}_{adv}$;
4 Experiments and Results
Figure 3: We present two examples of failed attacks. GCG-V produces benign content after affirmatively responding to malicious user requests, while VAJM generates harmful content without strictly adhering to user instructions.
4.1 Experimental Setup
4.1.1 Datasets
For the corpus of harmful sentences $S:=\{s_{i}\}_{i=1}^{m}$, we utilize the same derogatory corpus as VAJM (Qi et al., 2023), which includes 66 toxic sentences against gender, race, and humanity. Regarding the corpus of goal-target pairs $D:=\{g_{i},t_{i}\}_{i=1}^{n}$, we randomly select 66 samples from Advbench’s (Zou et al., 2023) harmful behaviors to serve as optimization samples. Advbench’s harmful behaviors comprises 520 harmful instructions paired with affirmative responses, primarily focusing on suggestions or guidance that are dangerous or illegal, but also encompassing other types of content. To rigorously assess the performance of various methods in jailbreak attacks, we first prompt the VLMs with 100 instructions randomly selected from the held-out subset of Advbench’s (Zou et al., 2023) harmful behaviors, which mainly focus on eliciting dangerous or illegal suggestions. We then extend our assessment by employing the VAJM (Qi et al., 2023) evaluation set, which comprises 40 detrimental instructions across four distinct categories: Identity Attack, Disinformation, Violence/Crime, and Malicious Behaviors toward Humanity (X-risk). To further broaden our evaluation scope, we also employ an automated testing procedure using the RealToxicityPrompts benchmark (Gehman et al., 2020). Specifically, we focus on a particularly challenging subset of RealToxicityPrompts (Schick et al., 2021; Mehrabi et al., 2022; Qi et al., 2023), consisting of 1,225 text prompts for triggering toxic continuations, which enables us to comprehensively evaluate the extent of harmful content produced by attacks.
4.1.2 Metrics
Here we describe the metrics used in our evaluations. Attack Success Rate (ASR). Given the tendency of keyword-based automated assessments to overinflate ASRs, we implement strict manual evaluation on both the Advbench’s (Zou et al., 2023) harmful behaviors and the VAJM (Qi et al., 2023) evaluation set. Attacks that generate relevant and useful harmful output are defined as successful, with others considered failures. Toxicity Rate. To measure the toxicity of continuations generated on the RealToxicityPrompts benchmark (Gehman et al., 2020), we utilize the Perspective API and the Detoxify classifier (Hanu & Unitary team, 2020), both of which can calculate toxicity scores for a set of six toxic attributes. Scores range from 0 (least toxic) to 1 (most toxic). For each attribute, we calculate the proportion of texts whose toxicity scores exceed a 0.5 threshold.
4.1.3 Implementation Details
To evaluate our attack methodology, we employ the Vicuna-13B version of MiniGPT-4 (Zhu et al., 2023). It is built on the frozen Vicuna-13B (Chiang et al., 2023) backbone model, an aligned LLM derived from LLaMA (Touvron et al., 2023). For the experimental setup, we adopt the configuration utilized in VAJM (Qi et al., 2023), specifying a batch size $b$ of 8, a step size $\alpha$ of 1, and an iteration count $N_{1}$ of 5000. Furthermore, we set the image-text attack ratio $r$ to 10, the length of the adversarial text suffix to 20 tokens, the number of candidates $n$ to 250, and the iteration count $N_{2}$ to 2000 by default. We use a single A100 GPU with 80GB of memory in all experiments.
4.2 Comparison With Unimodal Attacks
In this subsection, we compare the proposed UMK with state-of-the-art unimodal jailbreak attacks under different evaluation settings. Greedy Coordinate Gradient (GCG) (Zou et al., 2023) is a universal text-based attack devised for LLMs, which optimizes an adversarial text suffix to enhance the model’s propensity for generating affirmative responses. VAJM (Qi et al., 2023) targets VLMs with an image-based attack methodology, aiming to universally jailbreak VLMs by maximizing the probability of generating harmful sentences from a few-shot corpus. We reproduce GCG (Zou et al., 2023) and VAJM (Qi et al., 2023) with the official code and implement a visual version of GCG for a more comprehensive comparison.
Table 1: Comparison of Train ASR (%) and Test ASR (%) on Advbench’s (Zou et al., 2023) harmful behaviors. GCG-V refers to the visual version of GCG.
Method Train ASR (%) Test ASR (%)
Without Attack / 37.0
GCG (Zou et al., 2023) 68.2 50.0
VAJM (Qi et al., 2023) / 64.0
GCG-V 89.4 83.0
UMK(Ours) 100.0 96.0
In Table 1, the train ASR (%) and test ASR (%) of our proposed method are compared with baseline unimodal attacks on 66 training samples from Advbench’s (Zou et al., 2023) harmful behaviors and 100 test samples from the held-out set. GCG demonstrates markedly lower performance compared to visual attacks, with a test ASR of only 50%. VAJM (Qi et al., 2023) is trained on a corpus of harmful sentences $S:=\{s_{i}\}_{i=1}^{m}$, rather than being trained to generate affirmative responses to user queries. Consequently, we do not report its train ASR for Advbench’s harmful behaviors (Zou et al., 2023). Training on augmenting toxicity compromises VAJM’s capacity for instruction adherence. This approach leads to the generation of responses with high toxicity levels that fail to strictly follow user commands, culminating in a test ASR of merely 64%. GCG-V is the visual version of GCG, optimizing adversarial images to elicit affirmative responses. Similar to VAJM, GCG-V operates in the image domain. However, its training on Advbench’s harmful behaviors specifically enhances its ability to generate dangerous or illegal suggestions. Nonetheless, it still encounters conflicts between benign alignment and task completion, achieving a test ASR of 83%. Figure 3 showcases examples of failed attacks for each method. In comparison, the proposed UMK has addressed the aforementioned two issues and further improved performance through multimodal attacks, achieving an impressive 100% train ASR and 96% test ASR.
Table 2: Comparison of average ASR (%) measured on the VAJM (Qi et al., 2023) evaluation set across 4 categories of harmful instructions.
ASR(%) Identity Attack Disinfo Violence/Crime X-risk Without Attack 30.8 53.3 57.3 33.3 GCG (Zou et al., 2023) 49.2 48.9 57.3 40.0 VAJM (Qi et al., 2023) 81.5 82.2 85.3 60.0 GCG-V 66.2 64.4 84.0 6.7 UMK (Ours) 87.7 95.6 98.7 46.7
Table 3: Percentages (%) of outputs that display speciffc toxic attributes evaluated by the Perspective API and Detoxify classifier (Hanu & Unitary team, 2020). These outputs are generated on the ‘challenging’ subset from RealToxicityPrompts benchmark (Gehman et al., 2020). ‘Any*’ indicates the text exhibits at least one of the six toxic attributes.
Perspective API (%)
(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
Without Attack 1.60 25.23 1.18 12.62 1.60 29.18 34.31
VAJM (Qi et al., 2023) 18.05 43.44 7.69 14.70 6.35 62.66 67.17
GCG-V 6.42 57.05 2.09 21.02 3.67 64.22 71.98
UMK (Ours) 7.34 62.80 1.67 22.77 2.75 72.56 76.98
Detoxify (%)
(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
Without Attack 1.18 21.63 0.34 10.10 1.01 28.03 28.62
VAJM (Qi et al., 2023) 10.79 44.06 2.93 33.70 3.68 61.79 61.96
GCG-V 2.92 48.50 1.09 21.37 1.50 54.76 56.34
UMK (Ours) 3.92 58.51 0.92 28.46 1.00 67.86 68.70
For the VAJM (Qi et al., 2023) evaluation set, we employ nucleus sampling (Holtzman et al., 2019) with $p=0.9$ and temperature $=1$, generating five independent outputs for each instruction. The average attack success rate for harmful instructions in each category is reported in Table 2. The text-based attack, GCG (Zou et al., 2023), still results in poor performance. In contrast, VAJM (Qi et al., 2023) consistently outperforms GCG-V across all four categories of harmful instructions. This discrepancy in performance can be attributed to the differing training objectives of the two methods. VAJM is specifically designed to generate toxic sentences against gender, race, and humanity, leading to significantly better performance than GCG-V in categories such as identity attack, disinformation, and x-risk, achieving attack success rates of 81.5%, 82.2%, and 60.0%, respectively. Due to the Advbench’s (Zou et al., 2023) harmful behaviors being more closely aligned with the theme of violence/crime, GCG-V achieves a higher success rate in this category, while it reaches a notably low success rate of 6.7% in x-risk. Compared to others, our method achieves the best overall attack success rate. Notably, it even reaches an astonishing success rate of 98.7% in the category of violence/crime. However, it slightly falls short of VAJM’s performance in the x-risk category. This discrepancy is because the majority of harmful instructions in the x-risk category are formulated as questions, which differ from the declarative or directive formats used in the optimization samples for our adversarial attacks.
Table 4: Comparison of Train ASR (%) and Test ASR (%) for ablation studies on Advbench’s (Zou et al., 2023) harmful behaviors.
Method Train ASR (%) Test ASR (%)
Dual Objectives+Unimodal Attack 92.4 89.0
Single Objective+Multimodal Attack 95.5 92.0
Dual Objectives+Multimodal Attack 100.0 96.0
To further broaden our evaluation, we utilize the challenging subset of RealToxicityPrompts benchmark (Gehman et al., 2020), which includes 1,225 text prompts designed to trigger toxic continuations. Given that this is a text continuation task, we exclude the GCG method (Zou et al., 2023), which focuses on generating text suffixes, from our comparison. Similarly, for our approach, we utilize only the adversarial image prefix. We pair each adversarial image with text prompts from the dataset as input, and then measure the toxicity of the outputs using both the Perspective API and the Detoxify classifier [31], each capable of calculating toxicity scores for a set of six toxic attributes. For each attribute, we calculate the ratio of the generated texts with scores exceed the threshold of 0.5. As shown in Table 3, VAJM (Qi et al., 2023) and GCG-V display different performances across two evaluation methods. Under the Perspective API, GCG-V achieves a 71.98% toxicity rate, higher than VAJM’s 67.17%. Conversely, with the Detoxify classifier (Hanu & Unitary team, 2020), VAJM records a 61.96% rate, exceeding GCG-V’s 56.34%. Although VAJM achieves results superior to those of our method in certain categories such as Identity Attack, this is attributable to its optimization objective being primarily focused on generating harmful statements against gender, race, and humanity. In contrast, our method’s optimization objectives are more comprehensive. Although our method only utilizes adversarial image prefix in this experiment, it achieves overall best results, recording Any* toxicity rates of 76.98% and 68.70% under the Perspective API and Detoxify classifier tests, respectively. These results significantly outperform the previous best rates of 71.98% from GCG-V and 61.96% from VAJM. This demonstrates the effectiveness of our dual optimization objective strategy: even though the second phase is aimed at guiding the model to generate affirmative responses, it successfully enhances the toxic semantics of the adversarial image. Moreover, the joint attack of the text-image multimodalities may have uncovered a more optimal solution space for the adversarial image.
4.3 Ablation Studies
In this subsection, we conduct ablation studies on two key designs of the proposed attack strategy, i.e., the dual optimization objectives and the text-image multimodal attack. The unimodal attack method, which employs dual optimization objectives to guide the model towards affirmative responses after injecting toxic semantics within the image domain, achieves an 88% test ASR. The multimodal attack method, leveraging a single optimization objective to guide the model towards affirmative responses, achieves a 92% test ASR. Unimodal Attack with dual optimization objectives, compared to GCG-V shown in Table 1, results in a 6% increase in test ASR, demonstrating that the dual optimization objectives help induce the model to generate affirmative responses with higher toxicity. Moreover, the multimodal attack strategy utilizing a single optimization objective yields better results than the unimodal attack with dual optimization objectives, underscoring the effectiveness of multimodal attacks. However, both methods are surpassed by our proposed multimodal attack strategy with dual optimization objectives. This strategy begins by injecting toxic semantics into the adversarial image, then jointly optimizes both modalities to guide the model towards affirmative responses, achieving a 96% test ASR.
4.4 Understanding the Role of Text Attack in Multimodal Strategy
Figure 4: Comparative Analysis of Loss Before and After Text Attack. The X-axis represents ‘Steps’, while the Y-axis denotes ‘Loss’.
Figure 5: Overview of the image-image attack strategy. We adopt the same dual optimization objectives as used in the text-image attack.
Figure 6: Challenging examples of our proposed multimodal attack strategy that circumvents MiniGPT-4’s (Zhu et al., 2023) safety mechanisms, showcasing generated content that promotes harmful behaviors and makes contentious political comparisons.
In this subsection, we explore the role of text in the text-image multimodal attack strategy. Figure 4 illustrates the variance in batch loss before and after text attack throughout the execution of our attack methodology. For ease of analysis, we employ linear regression models to fit the curve. Observations at different steps indicate a significant reduction in batch loss after the text attack. As the number of steps increases, an overall downward trend in loss can be observed, demonstrating the efficacy of our method. Notably, even in the later stages of the attack, the reduction in batch loss before and after the text attack remains evident, emphasizing the pivotal contribution of text in amplifying the effectiveness of our strategy.
To better understand the role of text in our proposed text-image multimodal attack strategy, we introduce an image-image attack for comparative analysis. As illustrated in Figure 5, during the attack, we input two adversarial images and optimized them simultaneously. In the case of MiniGPT-4 (Zhu et al., 2023), the adversarial images we input each occupies 32 tokens, while the adversarial text suffix proposed in our attack utilizes only 20 tokens. Given that the textual space being discrete and denser compared to the visual space, adversarial attacks in the textual domain are generally more demanding. Intuitively, one might anticipate that image-image attacks, which are optimized over a larger volume of tokens, would yield superior results. However, this is not the case. In Table 5, we report the average ASR for both the image-image attack and the proposed text-image attack on the VAJM (Qi et al., 2023) evaluation set. It is evident that the efficacy of the image-image attack is significantly inferior to that of our attack. We observe that the image-image attacks are more likely to trigger the model to repeat affirmative responses twice or to cease output after the affirmative response. We believe this is because the optimization objectives for both adversarial images in the second phase are to maximize the probability of the model’s affirmative responses. In this context, the semantics of the two images tend to overshadow that of the text input. Consequently, during the testing phase, two adversarial images containing the same semantics are more likely to cause the model to duplicate affirmative reactions or only produce affirmative responses. On the other hand, this phenomenon can also be attributed to the training strategies employed by VLMs. Since VLMs are typically trained on image-text pairs, they tend to experience a performance degradation when encountering out-of-distribution inputs such as image-image-text trios.
4.5 Qualitative Analysis
In Figure 6, we present two challenging attack examples. Our text-image multimodal attack strategy can effectively circumvent the safety alignments of MiniGPT-4 (Zhu et al., 2023), generating content such as messages that encourage self-harm or suicide and comments that comparing politicians to Nazis.
Table 5: Comparative analysis of average ASR (%) between image-image and text-image attacks on the VAJM (Qi et al., 2023) evaluation set.
ASR(%) Identity Attack Disinfo Violence/Crime X-risk Image-Image Attack 78.5 68.9 85.3 6.7 Text-Image Attack 87.7 95.6 98.7 46.7
5 Conclusion
In this paper, we propose a text-image multimodal attack strategy with dual optimization objectives that can effectively jailbreak Large Vision-Language Models. Specifically, we first initialize the adversarial image prefix with random noise, then optimize it to generate harmful sentences without any text input, thereby infusing toxic semantic information into the image. We then introduce an adversarial text suffix and jointly optimize both the adversarial image prefix and the adversarial text suffix to generate affirmative responses to malicious user requests. By employing these two optimization objectives, we address the issues of insufficient toxicity in generated responses and the inability to adequately follow instructions. Through our text-image multimodal attack, we exploit a broader spectrum of attack surfaces exposed in VLMs, thereby enhancing the effectiveness of the attack. Experimental results indicate that our method significantly outperforms previous state-of-the-art unimodal attack approaches, achieving an attack success rate of 96% on MiniGPT-4 (Zhu et al., 2023). However, a limitation of the proposed method is its constrained transferability. We believe this is due to the varying model architectures, parameters, and even tokenizers among different VLMs. The proposed Universal Master Key (UMK), which presents itself in a form close to gibberish, carries semantic information that varies significantly across models, resulting in poor transferability. Enhancing this aspect of the proposed attack constitutes a significant direction for our future research.

### v2

@@
 White-box Multimodal Jailbreaks Against Large Vision-Language Models
-Abstract
+Abstract.
 Recent advancements in Large Vision-Language Models (VLMs) have underscored their superiority in various multimodal tasks.
@@
 Different from existing attacks, in this work we propose a more comprehensive strategy that jointly attacks both text and image modalities to exploit a broader spectrum of vulnerability within VLMs.
-Specifically, we propose a dual optimization objective aimed at guiding the model to generate affirmative responses with high toxicity.
+Specifically, we propose a dual optimization objective aimed at guiding the model to generate highly toxic affirmative responses.
 Our attack method begins by optimizing an adversarial image prefix from random noise to generate diverse harmful responses in the absence of text input, thus imbuing the image with toxic semantics.
@@
 When integrated into various malicious queries, UMK can circumvent the alignment defenses of VLMs and lead to the generation of objectionable content, known as jailbreaks.
-The experimental results demonstrate that our universal attack strategy can effectively jailbreak MiniGPT-4 with a 96% success rate, highlighting the vulnerability of VLMs and the urgent need for new alignment strategies.
-Disclaimer: This paper may contain content that is potentially disturbing and offensive.
-1 Introduction
+The experimental results demonstrate that our universal attack strategy can effectively jailbreak MiniGPT-4 with a 96% success rate, highlighting the fragility of VLMs and the exigency for new alignment strategies.
+Codes are available at https://github.com/roywang021/UMK.
+Disclaimer: This paper contains potentially disturbing and offensive content.
+Multimodal Jailbreak, White-box Jailbreak, Universal Adversarial Attacks, Vision-Language Models
+journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681092isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Artificial intelligence
+1. Introduction
+Figure 1. Example of Jailbreak Attack on MiniGPT-4 (Zhu et al., 2023). The proposed Universal Master key (UMK) helps arbitrary harmful queries bypass alignment constraints.
 The recent advancements in Large Vision-Language Models (VLMs) such as OpenAI’s GPT-4 (Achiam et al., 2023) and Google’s Flamingo (Alayrac et al., 2022) have received broad attention from academics, industry, and the general public.
-These developments have also garnered widespread attention from AI safety researchers, who are increasingly focused on assessing the adversarial robustness of multimodal models.
+These developments have also garnered widespread attention from AI safety researchers, who are increasingly focused on assessing the adversarial robustness of multimodal models (Schlarmann and Hein, 2023; Zhao et al., 2024).
 While VLMs are expected to exhibit the same alignment characteristics as Large Language Models (LLMs), providing helpful responses to user queries while rejecting requests that may cause harm, the integration of additional visual modality to base language models introduces novel vulnerabilities.
-Several methods (Carlini et al., 2024; Bagdasaryan et al., 2023; Bailey et al., 2023; Qi et al., 2023; Shayegani et al., 2023a) have successfully manipulated VLMs to produce harmful content by launching single-modality attacks on the continuous and high-dimensional visual modality.
+Several methods (Carlini et al., 2024; Bagdasaryan et al., 2023; Bailey et al., 2023; Qi et al., 2023; Shayegani et al., 2023a) have successfully manipulated VLMs to produce unexpected content by launching single-modality attacks on the continuous and high-dimensional visual modality.
 Conversely, research on purely text-based attacks is limited, attributed to a prevailing consensus regarding their ineffectiveness in breaching the defense mechanisms of well-aligned VLMs, due to the discrete and lower-dimensional nature of text (Carlini et al., 2024).
-Figure 1: Example of Jailbreak Attack on MiniGPT-4 (Zhu et al., 2023). The proposed Universal Master key (UMK) helps arbitrary harmful queries bypass alignment constraints.
-In contrast, this paper presents an initial effort in introducing a text-image multimodal attack strategy, aiming to uncover a wide range of intrinsic vulnerabilities within VLMs.
+In contrast, this paper presents an initial effort to introduce a text-image multimodal attack strategy, aiming to uncover a wide range of intrinsic vulnerabilities within VLMs.
 Furthermore, we have observed that while the previous unimodal universal attack strategy (Qi et al., 2023) increases response toxicity by perturbing images in the absence of text input, it also undermines the model’s adherence to instructions.
@@
 We expect that by attaching the UMK to arbitrary malicious user queries, it will circumvent the alignment defenses of VLMs and provoke the generation of objectionable content.
-To discover such a UMK, we start by initializing an adversarial image prefix from random noise and optimizing it without text input to maximize the model’s probability in generating harmful content, thereby imbuing the adversarial image prefix with toxic semantics.
+To discover such a UMK, we start by initializing an adversarial image prefix from random noise and optimizing it without text input to maximize the model’s probability of generating harmful content, thereby imbuing the adversarial image prefix with toxic semantics.
 Subsequently, we introduce an adversarial text suffix for joint optimization with the image prefix to maximize the probability of affirmative model responses.
@@
 The main contributions of our work can be summarized as:
-  • To the best of our knowledge, we are the first to introduce text-image multimodal adversarial attack against VLMs, systematically exploiting the vulnerabilities inherent in these models.
+  • To the best of our knowledge, we are the first to introduce text-image multimodal adversarial attacks against VLMs, systematically exploiting the vulnerabilities inherent in these models.
   • We propose a multimodal attack strategy with dual optimization objectives.
@@
   • Extensive experiments on benchmark datasets demonstrate that the proposed Universal Master Key (UMK) can universally jailbreak the VLMs with a remarkable success rate, surpassing existing unimodal methods.
-Figure 2: Overview of our multimodal attack strategy: The Universal Master Key (UMK) comprises an adversarial image prefix $X^{p}_{adv}$ and an adversarial text suffix $X^{s}_{adv}$.
+Figure 2.
+Overview of our multimodal attack strategy: The Universal Master Key (UMK) comprises an adversarial image prefix $X^{p}_{adv}$ and an adversarial text suffix $X^{s}_{adv}$.
 We first optimize $X^{p}_{adv}$ to maximize the generation probability of harmful content without text input to infuse toxic semantics.
 Subsequently, we concatenate the malicious user query with $X^{s}_{adv}$, and jointly optimize $X^{p}_{adv}$ and $X^{s}_{adv}$ to maximize the generation probability of affirmative responses, e.g., ‘Sure, here’s instruction for doing ******** (a bad thing)’.
-2 Related Work
-2.1 Large Vision-Language Models
+2. Related Work
+2.1. Large Vision-Language Models
 Large Vision-Language Models (VLMs) are vision-integrated Large Language Models (LLMs) that receive input in text and image formats and generate free-form text output for multimodal tasks.
@@
 In our work, we evaluate the robustness of VLMs against the proposed multimodal attack, revealing an alarming attack success rate of 96% on MiniGPT-4 (Zhu et al., 2023), emphasizing the urgent need for new alignment strategies to rectify this critical vulnerability.
-2.2 Attacks Against Multimodal Models
+2.2. Attacks Against Multimodal Models
 To attack multimodal models, Greshake et al.
 (Greshake et al., 2023) explored the effectiveness of manually injecting deceptive text into input images.
+Gong et al.
+(Gong et al., 2023) converted the harmful content into images through typography to bypass the safety alignment.
 In contrast, Other studies have proposed more sophisticated image-domain adversarial attack methods.
@@
 (Qi et al., 2023) adopts principles similar to Bagdasaryan et al.
-(Bagdasaryan et al., 2023), aiming to find an universal adversarial visual input.
+(Bagdasaryan et al., 2023), aiming to find a universal adversarial visual input.
 Specifically, the attack no longer focuses on specific output sentences but tries to maximize the probability of generating derogatory output from a corpus containing 66 harmful sample sentences.
@@
 (Qi et al., 2023) did not provide a name for their method, we designate it as the Visual Adversarial Jailbreak Method (VAJM) for easier reference in subsequent sections.
-Shayegani et al.
-(Shayegani et al., 2023a) proposed attacking publicly available visual encoders such as CLIP (Radford et al., 2021) used in the multimodal models, thereby eliminating the need for complete white-box access.
-While the aforementioned approaches have demonstrated impressive results, they mainly focus on exploring the adversarial robustness under the unimodal attacks, based on the consensus that attacking in the continuous image space is more effective than attacking in the discrete token space.
-However, this focus may lead to underutilization of the full attack surfaces available in multimodal models.
+Concurrent and independent from our work, Niu et al.
+(Niu et al., 2024) adopted the same optimization objective as our second stage to generate adversarial images on VLMs, then endeavored to transform them into adversarial text suffixes for LLMs.
+While the aforementioned approaches have demonstrated impressive results, they mainly focus on exploring the adversarial robustness under the unimodal attacks.
 In contrast, our proposed method explores a broader range of vulnerabilities inherent in VLMs through attacks on text-image multimodalities.
-Moreover, the universal adversarial attack method proposed by Qi et al.
-(Qi et al., 2023) has been found to compromise the model’s adherence to user instructions.
-In comparison, our method employs dual optimization objectives aimed at guiding the model to generate affirmative responses with high toxicity, effectively alleviating this issue.
-3 Methodology
-3.1 Threat Model
+Additionally, the universal adversarial attack method proposed by Qi et al.
+(Qi et al., 2023) has been found to compromise the model’s adherence to user instructions and the single optimization objective of producing affirmative responses, as employed in concurrent work (Niu et al., 2024), could result in insufficient toxicity in the model’s output.
+In comparison, our method utilizes dual optimization objectives aimed at guiding the model to generate highly toxic affirmative responses, effectively alleviating this concern.
+3. Methodology
+3.1. Threat Model
 Attack Goals.
 We consider single-turn conversations between a malicious user and a VLM chatbot.
-The attacker attempts to trigger a number of harmful behaviors by circumventing VLM’s security mechanisms, e.g., generating unethical content or dangerous instructions restricted by system prompts or RHLF alignment technique.
+The attacker attempts to trigger a number of harmful behaviors by circumventing VLM’s security mechanisms, e.g., generating unethical content or dangerous instructions restricted by system prompts or the RHLF alignment technique.
 Adversary Capabilities.
@@
 These models are sufficient to provide attackers with the insights necessary to generate malicious content, such as detailed instructions for making a bomb or materials that promote gender discrimination.
-3.2 Proposed Attack
-3.2.1 Formalization
+3.2. Proposed Attack
+3.2.1. Formalization
 For simplicity, we omit the implementation details of converting raw input to feature embeddings.
@@
 ‘||||| |’ denotes string concatenation.
-3.2.2 Methodology Intuition
+3.2.2. Methodology Intuition
 VAJM (Qi et al., 2023) generates universal adversarial images by guiding the model to produce harmful content in the absence of text input, but this approach undermines the model’s ability to faithfully follow user instructions.
@@
 The models still face conflicting objectives between task completion and benign alignment.
-Inspired by VAJM (Qi et al., 2023) and GCG (Zou et al., 2023), we propose a dual optimization objective that combines the advantages of both methods to maximize the probability of the model generating affirmative responses with high toxicity.
+Inspired by VAJM (Qi et al., 2023) and GCG (Zou et al., 2023), we propose a dual optimization objective that combines the advantages of both methods to maximize the probability of the model generating highly toxic affirmative responses.
 Specifically, we begin by compromising the model’s ethical alignment through the infusion of toxic semantics into the adversarial image prefix.
 Subsequently, we jointly optimize the adversarial image prefix and adversarial text suffix to encourage the model to generate affirmative responses, with the expectation that the model will accomplish tasks in a state conducive to producing harmful content.
-3.2.3 Embedding Toxic Semantics into Adversarial Image Prefix
+3.2.3. Embedding Toxic Semantics into Adversarial Image Prefix
 Inspired by (Qi et al., 2023), we establish a corpus containing several few-shot examples of harmful sentences $S:=\{s_{i}\}_{i=1}^{m}$.
@@
 This optimization problem can be efficiently solved using prevalent techniques in image adversarial attacks, such as Projected Gradient Descent (PGD) (Madry et al., 2017).
-3.2.4 Text-Image Multimodal Optimization for Maximizing Affirmative Response Probability
+3.2.4. Text-Image Multimodal Optimization for Maximizing Affirmative Response Probability
 To maximize the probability of generating affirmative responses, we further introduce an adversarial text suffix $X^{s}_{adv}$ in conjunction with the adversarial image prefix $X^{p}_{adv}$, which is imbued with toxic semantics.
@@
 The shared feature space between the image embeddings and the text token embeddings in VLMs facilitates concurrent updates of both $X^{p}_{adv}$ and $X^{s}_{adv}$ through a single backpropagation pass.
-This mechanism is analogous to the simultaneous update of distinct pixels in a single image in conventional image adversarial attack.
+This mechanism is analogous to the simultaneous update of distinct pixels in a single image in a conventional image adversarial attack.
 We employ Projected Gradient Descent (PGD) (Madry et al., 2017) to update the adversarial image prefix $X^{p}_{adv}$, and Greedy Coordinate Gradient (GCG) (Zou et al., 2023) to update the adversarial text suffix $X^{s}_{adv}$.
 GCG is the current state-of-the-art text-based attack strategy against LLMs.
-It utilizes the gradients with respect to the one-hot token indicators to identify a set of promising replacement candidates at each token position, and then chooses the substitution that results in the maximum loss reduction via a forward pass.
+It utilizes the gradients with respect to the one-hot token indicators to identify a set of promising replacement candidates at each token position and then chooses the substitution that results in the maximum loss reduction via a forward pass.
 The overview of our attack is illustrated in Figure 2, while the detailed strategy is summarized in Algorithm 1.
 Algorithm 1 Multimodal Attack Strategy with Dual Optimization Objectives
-0:  VLM model $f_{\theta}$, harmful sentences corpus $S:=\{s_{i}\}_{i=1}^{m}$, goal-target pairs corpus $D:=\{g_{i},t_{i}\}_{i=1}^{n}$, image-text attack ratio $r$, batch size $b$, step size $\alpha$, number of candidates $n$, iteration counts $N_{1},N_{2}$;
-1:  Initialize $X^{p}_{adv}$ with random noise;
-2:  for $k=1,...,N_{1}$ do
-3:     Select $b$ samples from the corpus $S$, forming $S_{k}:=\{s^{\prime}_{i}\}_{i=1}^{b}$;
-4:     $\mathcal{L}(X^{p}_{adv},S_{k})=\sum_{i=1}^{b}-\log\left(p(s^{\prime}_{i}\mid X^{p}_{adv},\,\emptyset)\right)$;
-5:     $X^{p}_{adv}=\operatorname{clip}\left(X^{p}_{adv}+\alpha\operatorname{sign}\left(\nabla_{x^{p}}\mathcal{L}(X^{p}_{adv},S_{k})\right)\right)$;
-6:  end for
-7:  Initialize $X^{s}_{adv}$ using a predefined strategy;
-8:  for $j=1,...,N_{2}$ do
-9:     Select $b$ samples from the corpus $D$, forming $D_{j}:=\{g^{\prime}_{i},t^{\prime}_{i}\}_{i=1}^{b}$;
-10:     $\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})=\sum_{i=1}^{b}-\log\left(p(t^{\prime}_{i}\mid X^{p}_{adv},\,g^{\prime}_{i}||X^{s}_{adv})\right)$;
-11:     $X^{p}_{adv}=\operatorname{clip}\left(X^{p}_{adv}+\alpha\operatorname{sign}\left(\nabla_{x^{p}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})\right)\right)$;
-12:     if $j\%r==0$ then
-13:        Compute coordinate gradient $\nabla_{x^{s}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})$;
-14:        Obtain a set of candidate suffixes with single-token substitution $\{X^{s}_{adv}\}^{n}$;
-15:        $X^{s}_{adv}=\underset{\{X^{s}_{adv}\}^{n}}{\operatorname{argmin}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})$;
-16:     end if
-17:  end for
-18:  Obtain the Universal Master Key (UMK) composed of $X^{p}_{adv}$ and $X^{s}_{adv}$;
-4 Experiments and Results
-Figure 3: We present two examples of failed attacks.
+1:VLM model $f_{\theta}$, harmful sentences corpus $S:=\{s_{i}\}_{i=1}^{m}$, goal-target pairs corpus $D:=\{g_{i},t_{i}\}_{i=1}^{n}$, image-text attack ratio $r$, batch size $b$, step size $\alpha$, number of candidates $n$, iteration counts $N_{1},N_{2}$;
+2:Initialize $X^{p}_{adv}$ with random noise;
+3:for $k=1,...,N_{1}$ do
+4:     Select $b$ samples from the corpus $S$, forming $S_{k}:=\{s^{\prime}_{i}\}_{i=1}^{b}$;
+5:     $\mathcal{L}(X^{p}_{adv},S_{k})=\sum_{i=1}^{b}-\log\left(p(s^{\prime}_{i}\mid X^{p}_{adv},\,\emptyset)\right)$;
+6:     $X^{p}_{adv}=\operatorname{clip}\left(X^{p}_{adv}+\alpha\operatorname{sign}\left(\nabla_{x^{p}}\mathcal{L}(X^{p}_{adv},S_{k})\right)\right)$;
+7:end for
+8:Initialize $X^{s}_{adv}$ using a predefined strategy;
+9:for $j=1,...,N_{2}$ do
+10:     Select $b$ samples from the corpus $D$, forming $D_{j}:=\{g^{\prime}_{i},t^{\prime}_{i}\}_{i=1}^{b}$;
+11:     $\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})=\sum_{i=1}^{b}-\log\left(p(t^{\prime}_{i}\mid X^{p}_{adv},\,g^{\prime}_{i}||X^{s}_{adv})\right)$;
+12:     $X^{p}_{adv}=\operatorname{clip}\left(X^{p}_{adv}+\alpha\operatorname{sign}\left(\nabla_{x^{p}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})\right)\right)$;
+13:     if $j\%r==0$ then
+14:         Compute coordinate gradient $\nabla_{x^{s}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})$;
+15:         Obtain a set of candidate suffixes with single-token substitution $\{X^{s}_{adv}\}^{n}$;
+16:         $X^{s}_{adv}=\underset{\{X^{s}_{adv}\}^{n}}{\operatorname{argmin}}\mathcal{L}(X^{p}_{adv},X^{s}_{adv},D_{j})$;
+17:     end if
+18:end for
+19:Obtain the Universal Master Key (UMK) composed of $X^{p}_{adv}$ and $X^{s}_{adv}$;
+4. Experiments and Results
+Figure 3.
+We present two examples of failed attacks.
 GCG-V produces benign content after affirmatively responding to malicious user requests, while VAJM generates harmful content without strictly adhering to user instructions.
-4.1 Experimental Setup
-4.1.1 Datasets
+4.1. Experimental Setup
+4.1.1. Datasets
 For the corpus of harmful sentences $S:=\{s_{i}\}_{i=1}^{m}$, we utilize the same derogatory corpus as VAJM (Qi et al., 2023), which includes 66 toxic sentences against gender, race, and humanity.
@@
 Specifically, we focus on a particularly challenging subset of RealToxicityPrompts (Schick et al., 2021; Mehrabi et al., 2022; Qi et al., 2023), consisting of 1,225 text prompts for triggering toxic continuations, which enables us to comprehensively evaluate the extent of harmful content produced by attacks.
-4.1.2 Metrics
+4.1.2. Metrics
 Here we describe the metrics used in our evaluations.
@@
 Toxicity Rate.
-To measure the toxicity of continuations generated on the RealToxicityPrompts benchmark (Gehman et al., 2020), we utilize the Perspective API and the Detoxify classifier (Hanu & Unitary team, 2020), both of which can calculate toxicity scores for a set of six toxic attributes.
+To measure the toxicity of continuations generated on the RealToxicityPrompts benchmark (Gehman et al., 2020), we utilize the Perspective API and the Detoxify classifier (Hanu and Unitary team, 2020), both of which can calculate toxicity scores for a set of six toxic attributes.
 Scores range from 0 (least toxic) to 1 (most toxic).
 For each attribute, we calculate the proportion of texts whose toxicity scores exceed a 0.5 threshold.
-4.1.3 Implementation Details
+4.1.3. Implementation Details
 To evaluate our attack methodology, we employ the Vicuna-13B version of MiniGPT-4 (Zhu et al., 2023).
@@
 Furthermore, we set the image-text attack ratio $r$ to 10, the length of the adversarial text suffix to 20 tokens, the number of candidates $n$ to 250, and the iteration count $N_{2}$ to 2000 by default.
-We use a single A100 GPU with 80GB of memory in all experiments.
-4.2 Comparison With Unimodal Attacks
+We use a single A100 GPU with 80GB of memory for all experiments.
+4.2. Comparison With Unimodal Attacks
 In this subsection, we compare the proposed UMK with state-of-the-art unimodal jailbreak attacks under different evaluation settings.
@@
 We reproduce GCG (Zou et al., 2023) and VAJM (Qi et al., 2023) with the official code and implement a visual version of GCG for a more comprehensive comparison.
-Table 1: Comparison of Train ASR (%) and Test ASR (%) on Advbench’s (Zou et al., 2023) harmful behaviors. GCG-V refers to the visual version of GCG.
+Table 1. Comparison of Train ASR (%) and Test ASR (%) on Advbench’s (Zou et al., 2023) harmful behaviors. GCG-V refers to the visual version of GCG.
 Method Train ASR (%) Test ASR (%)
@@
 In comparison, the proposed UMK has addressed the aforementioned two issues and further improved performance through multimodal attacks, achieving an impressive 100% train ASR and 96% test ASR.
-Table 2: Comparison of average ASR (%) measured on the VAJM (Qi et al., 2023) evaluation set across 4 categories of harmful instructions.
-ASR(%) Identity Attack Disinfo Violence/Crime X-risk Without Attack 30.8 53.3 57.3 33.3 GCG (Zou et al., 2023) 49.2 48.9 57.3 40.0 VAJM (Qi et al., 2023) 81.5 82.2 85.3 60.0 GCG-V 66.2 64.4 84.0 6.7 UMK (Ours) 87.7 95.6 98.7 46.7
-Table 3: Percentages (%) of outputs that display speciffc toxic attributes evaluated by the Perspective API and Detoxify classifier (Hanu & Unitary team, 2020).
+Table 2. Comparison of average ASR (%) measured on the VAJM (Qi et al., 2023) evaluation set across 4 categories of harmful instructions.
+ASR(%) Identity Attack Disinfo Violence/Crime X-risk
+Without Attack 30.8 53.3 57.3 33.3
+GCG (Zou et al., 2023) 49.2 48.9 57.3 40.0
+VAJM (Qi et al., 2023) 81.5 82.2 85.3 60.0
+GCG-V 66.2 64.4 84.0 6.7
+UMK (Ours) 87.7 95.6 98.7 46.7
+Table 3.
+Percentages (%) of outputs that display specific toxic attributes evaluated by the Perspective API and Detoxify classifier (Hanu and Unitary team, 2020).
 These outputs are generated on the ‘challenging’ subset from RealToxicityPrompts benchmark (Gehman et al., 2020).
@@
 This discrepancy is because the majority of harmful instructions in the x-risk category are formulated as questions, which differ from the declarative or directive formats used in the optimization samples for our adversarial attacks.
-Table 4: Comparison of Train ASR (%) and Test ASR (%) for ablation studies on Advbench’s (Zou et al., 2023) harmful behaviors.
+Table 4. Comparison of Train ASR (%) and Test ASR (%) for ablation studies on Advbench’s (Zou et al., 2023) harmful behaviors.
 Method Train ASR (%) Test ASR (%)
@@
 Similarly, for our approach, we utilize only the adversarial image prefix.
-We pair each adversarial image with text prompts from the dataset as input, and then measure the toxicity of the outputs using both the Perspective API and the Detoxify classifier [31], each capable of calculating toxicity scores for a set of six toxic attributes.
-For each attribute, we calculate the ratio of the generated texts with scores exceed the threshold of 0.5.
+We pair each adversarial image with text prompts from the dataset as input and then measure the toxicity of the outputs using both the Perspective API and the Detoxify classifier [31], each capable of calculating toxicity scores for a set of six toxic attributes.
+For each attribute, we calculate the ratio of the generated texts with scores exceeding the threshold of 0.5.
 As shown in Table 3, VAJM (Qi et al., 2023) and GCG-V display different performances across two evaluation methods.
 Under the Perspective API, GCG-V achieves a 71.98% toxicity rate, higher than VAJM’s 67.17%.
-Conversely, with the Detoxify classifier (Hanu & Unitary team, 2020), VAJM records a 61.96% rate, exceeding GCG-V’s 56.34%.
-Although VAJM achieves results superior to those of our method in certain categories such as Identity Attack, this is attributable to its optimization objective being primarily focused on generating harmful statements against gender, race, and humanity.
+Conversely, with the Detoxify classifier (Hanu and Unitary team, 2020), VAJM records a 61.96% rate, exceeding GCG-V’s 56.34%.
+Although VAJM achieves results superior to those of our method in certain categories such as Identity Attacks, this is attributable to its optimization objective being primarily focused on generating harmful statements against gender, race, and humanity.
 In contrast, our method’s optimization objectives are more comprehensive.
-Although our method only utilizes adversarial image prefix in this experiment, it achieves overall best results, recording Any* toxicity rates of 76.98% and 68.70% under the Perspective API and Detoxify classifier tests, respectively.
+Although our method only utilizes adversarial image prefixes in this experiment, it achieves overall best results, recording Any* toxicity rates of 76.98% and 68.70% under the Perspective API and Detoxify classifier tests, respectively.
 These results significantly outperform the previous best rates of 71.98% from GCG-V and 61.96% from VAJM.
@@
 Moreover, the joint attack of the text-image multimodalities may have uncovered a more optimal solution space for the adversarial image.
-4.3 Ablation Studies
+Figure 4. Comparative Analysis of Loss Before and After Text Attack. The X-axis represents ‘Steps’, while the Y-axis denotes ‘Loss’.
+4.3. Ablation Studies
 In this subsection, we conduct ablation studies on two key designs of the proposed attack strategy, i.e., the dual optimization objectives and the text-image multimodal attack.
@@
 This strategy begins by injecting toxic semantics into the adversarial image, then jointly optimizes both modalities to guide the model towards affirmative responses, achieving a 96% test ASR.
-4.4 Understanding the Role of Text Attack in Multimodal Strategy
-Figure 4: Comparative Analysis of Loss Before and After Text Attack. The X-axis represents ‘Steps’, while the Y-axis denotes ‘Loss’.
-Figure 5: Overview of the image-image attack strategy. We adopt the same dual optimization objectives as used in the text-image attack.
-Figure 6: Challenging examples of our proposed multimodal attack strategy that circumvents MiniGPT-4’s (Zhu et al., 2023) safety mechanisms, showcasing generated content that promotes harmful behaviors and makes contentious political comparisons.
+4.4. Understanding the Role of Text Attack in Multimodal Strategy
+Figure 5. Overview of the image-image attack strategy. We adopt the same dual optimization objectives as used in the text-image attack.
+Figure 6.
+Examples of challenging attacks demonstrating our text-image multimodal strategy.
+The approach effectively bypasses MiniGPT-4’s (Zhu et al., 2023) safety mechanisms, generating harmful content including messages promoting self-harm or suicide and derogatory comparisons of politicians to Nazis.
 In this subsection, we explore the role of text in the text-image multimodal attack strategy.
-Figure 4 illustrates the variance in batch loss before and after text attack throughout the execution of our attack methodology.
+Figure 4 illustrates the variance in batch loss before and after the text attack throughout the execution of our attack methodology.
 For ease of analysis, we employ linear regression models to fit the curve.
@@
 Notably, even in the later stages of the attack, the reduction in batch loss before and after the text attack remains evident, emphasizing the pivotal contribution of text in amplifying the effectiveness of our strategy.
+Table 5. Comparative analysis of average ASR (%) between image-image and text-image attacks on the VAJM (Qi et al., 2023) evaluation set.
+ASR(%) Identity Attack Disinfo Violence/Crime X-risk
+Image-Image Attack 78.5 68.9 85.3 6.7
+Text-Image Attack 87.7 95.6 98.7 46.7
 To better understand the role of text in our proposed text-image multimodal attack strategy, we introduce an image-image attack for comparative analysis.
@@
 In the case of MiniGPT-4 (Zhu et al., 2023), the adversarial images we input each occupies 32 tokens, while the adversarial text suffix proposed in our attack utilizes only 20 tokens.
-Given that the textual space being discrete and denser compared to the visual space, adversarial attacks in the textual domain are generally more demanding.
+Given that the textual space is discrete and denser compared to the visual space, adversarial attacks in the textual domain are generally more demanding.
 Intuitively, one might anticipate that image-image attacks, which are optimized over a larger volume of tokens, would yield superior results.
@@
 On the other hand, this phenomenon can also be attributed to the training strategies employed by VLMs.
-Since VLMs are typically trained on image-text pairs, they tend to experience a performance degradation when encountering out-of-distribution inputs such as image-image-text trios.
-4.5 Qualitative Analysis
-In Figure 6, we present two challenging attack examples.
-Our text-image multimodal attack strategy can effectively circumvent the safety alignments of MiniGPT-4 (Zhu et al., 2023), generating content such as messages that encourage self-harm or suicide and comments that comparing politicians to Nazis.
-Table 5: Comparative analysis of average ASR (%) between image-image and text-image attacks on the VAJM (Qi et al., 2023) evaluation set.
-ASR(%) Identity Attack Disinfo Violence/Crime X-risk Image-Image Attack 78.5 68.9 85.3 6.7 Text-Image Attack 87.7 95.6 98.7 46.7
-5 Conclusion
+Since VLMs are typically trained on image-text pairs, they tend to experience performance degradation when encountering out-of-distribution inputs such as image-image-text trios.
+5. Conclusion
 In this paper, we propose a text-image multimodal attack strategy with dual optimization objectives that can effectively jailbreak Large Vision-Language Models.
@@
 Enhancing this aspect of the proposed attack constitutes a significant direction for our future research.
+Acknowledgements.
+This work was supported by the National Key R&D Program of China (Grant No. 2022ZD0160103) and the National Natural Science Foundation of China (Grant No. 62276067).